A class survey showed that 7 out of 25 students plan to travel over summer vacation. How many students out of the 200 students in the school would you expect to be traveling in the summer?
Answer by stanbon(75887)

A class survey showed that 7 out of 25 students plan to travel over summer vacation. How many students out of the 200 students in the school would you expect to be traveling in the summer?
-----------------------------
Use a proportion:
x/200 = 7/25
----
x = 200*7/25
---
x = 8*7 = 56
==================
Cheers,
Stan H.